# VeOmni: Scaling Any Modality Model Training with Model-Centric Distributed Recipe Zoo

### v2

@@ B.4 Parallel State of VeOmni
B.5 FSDP and parallelized
Listing LABEL:lst:veomni-build-parallelize-model illustrates the usage of the build_parallelize_model API, which applies Fully Sharded Data Parallelism (FSDP) and a user-defined parallel plan to the target model. This interface abstracts away low-level distributed training details, such as parameter sharding, precision configuration, and gradient checkpointing. As a result, users can enable efficient large-scale distributed training with minimal code changes, achieving scalability without compromising usability.
Listing 6: build_parallelize_model API
@@ B.5 FSDP and parallelized
B.6 Ulysess flash attention forward Function
Listing LABEL:lst:veomni-ulysess-flash-attention-forward presents the implementation of flash_attention_forward in VeOmni, which extends HuggingFace’s native FlashAttention interface to support Ulysses-style sequence parallelism. When sequence parallelism is enabled via the global parallel state, the input query, key, and value tensors are first transformed from sequence-sharded to head-sharded layouts using all-to-all collective operations. After performing attention computation through the standard _flash_attention_forward function, the output is transformed back from head-sharded to sequence-sharded format. This function maintains modular compatibility with HuggingFace models, requiring no changes to attention logic, and enables seamless integration of sequence parallelism with FlashAttention.
Listing 7: Ulysess flash attention forward Function
@@ B.6 Ulysess flash attention forward Function
B.7 Expert Parallel Plan and Expert Parallel Function
Listing LABEL:lst:veomni-ep-plan demonstrates the definition of an expert parallel (EP) plan using VeOmni’s ParallelPlan interface. This declarative design allows users to specify which modules (e.g., expert projection layers in an MoE block) should be sharded across expert devices. The syntax follows a wildcard-matching convention similar to PyTorch module names, and leverages DTensor-compatible placement primitives like Shard(0). When passed to build_parallelize_model(), the plan is automatically applied to partition the target model along expert dimensions. This flexible mechanism not only simplifies expert parallelism but also enables clean integration with other parallelism strategies such as tensor parallelism and FSDP.
Listing 8: ParallelPlan API

### v3

@@ B.4 Parallel State Management
B.5 Omni-Modal Model Parallelization
Listing LABEL:lst:veomni-build-parallelize-model illustrates the usage of the build_parallelize_model API, which applies fully sharded data parallel (FSDP) and a customized parallel plan to the target model. This interface abstracts away low-level distributed training details, such as parameter sharding, precision configuration, and activation checkpointing, enabling efficient large-scale distributed training with minimal code changes and achieving scalability without compromising usability.
Listing 6: Omni-Modal Model Parallelization
@@ B.5 Omni-Modal Model Parallelization
B.6 Long-Context Attention
Listing LABEL:lst:veomni-ulysess-flash-attention-forward presents the implementation of flash_attention_forward in VeOmni, which extends HuggingFace’s native FlashAttention interface to support Ulysses-style sequence parallelism [29]. When sequence parallelism is enabled, the query, key, and value tensors are first transformed from sequence-sharded to head-sharded layouts using all-to-all collective operations. After performing attention computation through the standard _flash_attention_forward function, the output is transformed back from head-sharded to sequence-sharded format. This function maintains modular compatibility with HuggingFace models, requiring no changes to attention logic, and enables seamless integration of sequence parallelism with FlashAttention [9].
Listing 7: Long-Context Attention
@@ B.6 Long-Context Attention
B.7 Expert Parallelism Implementation
Listing LABEL:lst:veomni-ep-plan demonstrates the definition of an expert parallelism (EP) plan using VeOmni’s ParallelPlan interface. This declarative design allows users to specify which modules (e.g., expert projection layers in an MoE block) should be sharded across expert parallelism groups. The syntax follows a wildcard-matching convention similar to PyTorch module names, and leverages DTensor-compatible placement primitives like Shard(0). When passed to build_parallelize_model(), the plan is automatically applied to partition the target model along expert dimensions. This flexible mechanism not only simplifies expert parallelism but also enables clean integration with other parallelism strategies such as tensor parallelism (TP) and FSDP.
Listing 8: Expert Parallelism Implementation